Fowler and Woods Enterprises is a publicly traded company that just paid a $2.00 per share dividend.  The company is expected to increase its dividend by 20% per year for the next two years.  After the second year, the dividend growth rate will be 5% per year for the next two years.  After the 4th year, dividends are expected to grow at a constant rate of 3% into the foreseeable future.  An analyst estimates that investors in the firm will require a 12% annual return.  Based on this information, what is the intrinsic value of the stock today?
Answer #1
Intrinsic value of the stock today is the PV of the expected dividends
when discounted at the required rate of return of 12%.
Year Dividend PVIF at 12% PV at 12%
0 $                 2.00
1 $                 2.40 0.89286 $               2.14
2 $                 2.88 0.79719 $               2.30
3 $                 3.02 0.71178 $               2.15
4 $                 3.18 0.63552 $               2.02
Sum of PV of dividends of years 1 to 4 $               8.61
Continuing value of dividends at t4 = 3.18*1.03/(0.12-0.03) = $              36.39
PV of continuing value = 36.39*0.63552 = $            23.13
Intrinsic value of the stock today = 8.61+23.13 = $            31.74